2 cars start from the same place and time, traveling in opposite directions, one averaging 65km p/h, the other 90km p/h. After how long will they be 450km apart ? : 2 cars start from the same place and time, traveling in opposite directions, one averaging 65km p/h, the other 90km p/h. After how long will they be 450km apart ?
Answer by aka042(26)

Hi-
Let's call the place the two cars start from the origin. If we look at the car travelling 65 km/h, then after a certain number of hours (lets call it h), the car will be 65km * h away from the origin. So after one hour, the car will be 65km * 1 = 65 km from the origin.
Similarly, the second car will be 90km * h away from the origin (and after one hour, it will be 90km * 1 = 90km from the origin).
The distance between the two cars is simply the sum of each car's distance from the origin. So after one hour, the two cars will be (65km * 1) + (90km * 1) = 155km from each other.
Let's generalize that by saying D = 65*h + 90 * h where D is the distance between the two cars after a certain number of hours h. We can simplify this equation to D = (65+90)*h which simplified further is D = 155*h.
Now, to solve the problem. You are given that the cars are 450 km away from each other, and need to find the time that has elapsed. So lets plug 450 into the equation and we get that 450 = 155*h. To solve, we simply divide each side by 450 (the coefficient of h, that way you will get h alone), and 450/155 = h, or h = 90/31 (about 2.90 hours).
Question 159376: 2 cars start from the same place and time, traveling in opposite directions, one averaging 65km p/h, the other 90km p/h. After how long will they be 450km apart ? : 2 cars start from the same place and time, traveling in opposite directions, one averaging 65km p/h, the other 90km p/h. After how long will they be 450km apart ?
Answer by Alan3354(1934)

2 cars start from the same place and time, traveling in opposite directions, one averaging 65km p/h, the other 90km p/h. After how long will they be 450km apart ?
-------------------
They're getting farther apart at the sum of their speeds.
65 + 90 = 155 kph
The time to get 450 km apart is 450/155
= 2.90322 hours
= 2 hours 54 minutes